A company is measuring the time it takes to complete an order. Twenty-five samples of size 6 resulted in an average of 80 minutes with an average range of 20 minutes. Find control limits for the appropriate control chart(s) that would be used to monitor this process.
Answer #1

Both x-bar and R-chart can be used

Mean = 80,

R-bar = 20.

From the table of control chart factors for a sample size of 6, A2= 0.483, D3= 0, and D4= 2.004.

Control limits for the x-bar chart are:

LCL = 80 - 0.483(20) = 70.34

UCL = 80 + 0.483(20) = 89.66

Control limits for the R-chart are: LCL = 0

UCL = 2.004(20) = 40.08